Five years ago I was thrice as old as my son and ten years later I shall be twice as old as my son. How old are we now?
Answer by lwsshak3(11628)

Five years ago I was thrice as old as my son and ten years later I shall be twice as old as my son. How old are we now?
let x=son's age now
5 yrs ago:
x-5=son's age
3(x-5)=parent's age
10 yrs later:
x+10=son's age
2(x+10)=parent's age
parent's age 10 yrs later=parent's age 5 yrs ago plus 15
2(x+10)=15+3(x-5)
2x+20=15+3x-15
x=20
3(x-5)+5=50
son's age now: 20
parent's age now: 50